OpenStudy (anonymous):

I REALLY need help on this question please. use the formula d=rt to find the value of the missing variable. d=135 feet, t=27 seconds r=___ft/sec

zepdrix (zepdrix):

\[\large d=rt\]\[\large \text{Given:} \quad d=135ft, \quad t=27s\]Plugging these values in for d and t gives us,\[\large d=rt \qquad \rightarrow \qquad 135ft=r\cdot 27s\]To solve for r, we'll divide both sides by 27.\[\large \frac{135}{27}\frac{ft}{s}=r\cdot \frac{27s}{27s}\]\[\large r=\frac{135}{27}\frac{ft}{s}\]Understand how that works? Punch it into your calculator if your teacher prefers a decimal answer.

OpenStudy (anonymous):

would r=135 ____ 27 be the correct answer or can it be simplified and if so how do i do it?

OpenStudy (noseboy908):

to simplify that, (if necessary), solve it: 135/27.

OpenStudy (anonymous):

so the answer would be r=135 over 27?

zepdrix (zepdrix):

Hmm it looks like 135 and 27 are both divisible by 9, so yes it could be simplified a little bit.

OpenStudy (noseboy908):

yes, but first you must simplify.

OpenStudy (anonymous):

so simplified would it be r= 5 over 1?

zepdrix (zepdrix):

Mmmm yah that sounds right :D

zepdrix (zepdrix):

Which can be simplified to 5

OpenStudy (anonymous):

but would I write the answer as r= 5 over 1? or r=5?

OpenStudy (noseboy908):

r=5. although, they technically mean the same thing.

OpenStudy (anonymous):

ok thank you very much I really appreciate it